Please help me to solve this equation: x^2-2x-3=0
Answer by Earlsdon(6294)

This quadratic equation can be solved by factoring.
x^2 - 2x - 3 = 0 Look for two integers whose product is -3 and whose sum is -2
Some possibilities: Since the product is negative, the two integers will have opposite signs (+, -) and since the sum is also negative, the larger (absolute value-wise) of the two integers will be negative.
1 X -3 = -3 and 1 + (-3) = -2 Bingo!
So the two integers are: 1 and -3
The factors of this quadratic equation are:
(x + 1)(x - 3) = 0 Apply the zero products principle:
x + 1 = 0; x = -1
x - 3 = 0; x = 3
The verification is left as an exercise for the student.